solve by using the square root property.
9x^2=14
and the other problem is
(x-5)^2=7

Answer by edjones(8007)

9x^2=14
3x=+-sqrt(14)
x=+-sqrt(14)/3
.
(x-5)^2=7
x-5=+-sqrt(7)
x=5+-sqrt(7)
.
Ed